OpenStudy (anonymous):

Evaluate the following expression using the values given. Find x + 2y if x = 3 and y = 4

OpenStudy (anonymous):

Sub in the values of x and y into the equation. x=3 and y=4 Equation: x +2y =3 +(2*4) =3+8 =11

OpenStudy (anonymous):

x+2y if x=3 and y=4 =3+2(4) =3+8 =11